OpenStudy (anonymous):

A 6.5×10^(−2) arrow hits the target at 25m/s and penetrates 3.8cm before stopping. What average force did the target exert on the arrow? 2 sig figs please

OpenStudy (anonymous):

find acceleration from 25^2 = 2*a*.038 F*delta t = m*delta V F= m *delta V / delta t = m*a